# Explainable Deep Convolutional Multi-Type Anomaly Detection

### v1

@@ 5.2 Multi-Class Anomaly Detection
We also evaluate MultiTypeFCDD in a multi-class anomaly detection setting using a single-channel anomaly output. This allows for fair benchmarking against state-of-the-art multi-class anomaly detection methods, as summarised in Table 5.2. While unsupervised methods show degradation in performance as the proportion of anomalous samples in the training dataset increases, weakly supervised approaches like ours, even with just image-level labels, benefit from additional anomalous samples and show improved performance. MultiTypeFCDD achieves consistently high AUROC scores across contamination levels, demonstrating strong anomaly detection capability and generalisation across object categories. Although its AUPRO scores, which indicate anomaly localisation performance, are lower than those of methods such as PatchCore or SoftPatch, this is mainly because these approaches train separate models for each object class, allowing more precise localisation, and the results are averaged. In contrast, unified frameworks like MultiTypeFCDD and UniAD utilise a single model trained across all object classes, which makes fine-grained localisation more challenging but substantially reduces the total number of parameters, requiring only a fraction of the memory and computational overhead while maintaining competitive overall performance.
Table 5: Multi-Class Anomaly Detection Performance (I-AUROC / P-AUPRO). A * indicates that the methods use a single unified model across all object classes.
Method $\alpha=0.1$ $\alpha=0.2$ $\alpha=0.4$ #Params
AUROC AUPRO AUROC AUPRO AUROC AUPRO (M)
PaDiM [defard2021padim] 81.9 86.4 80.1 86.5 77.0 86.1 68.9
CFlow [gudovskiy2022cflow] 83.9 90.7 79.6 90.7 78.0 90.2 30.4
PatchCore [roth2022towards] 90.4 94.2 89.5 93.9 88.1 92.4 68.9
SoftPatch [jiang2022softpatch] 90.9 92.9 90.5 92.9 89.3 92.5 68.9
SimpleNet [liu2023simplenet] 79.9 83.9 75.9 83.2 74.7 70.5 68.9
DeSTSeg [zhang2023destseg] 85.6 86.9 80.3 83.2 74.4 75.5 37.9
RD [deng2022anomaly] 95.1 87.6 94.9 83.2 94.7 81.5 68.9
UniAD[you2022unified] 84.2 87.7 82.8 87.3 80.1 86.6 7.6
MultiTypeFCDD* 94.1 75.2 95.5 75.4 96.4 79.9 5.5
@@ 5.3 Computational Performance
To emphasise the practical deployability of our approach, we benchmark its computational footprint against MultiADS [sadikaj2025multiads] presented in Table 6. MultiADS is a suitable baseline since it can segment multiple defect types similar to our method. However, it relies on image–text prompts to train a VLM, whereas our framework uses only image inputs and standard convolutional layers. Consequently, we do not perform a direct comparison under the multi-type anomaly detection setting.
It is to be noted that our model contains only 5.5 M parameters and occupies 19.7 MB of storage, over 80 times smaller than MultiADS, which requires 427.9 M parameters and 1.6 GB of memory. Furthermore, MultiTypeFCDD achieves significantly faster inference times, requiring only 69.9 ms per image on CPU and 1.7 ms on GPU, compared to 1824.3 ms (CPU) and 33.7 ms (GPU) for MultiADS. This comparison demonstrates that complex VLMs are not strictly necessary to achieve reliable multi-type anomaly detection. These findings align with our observations in Section 2.4, where large-scale VLMs are noted to be computationally and memory-intensive, restricting their usability in real-time or embedded systems [garcia2019estimation, marshall2025transformers, strubell2020energy]. In contrast, our lightweight architecture achieves competitive performance while substantially reducing memory and inference costs.
@@
Model Parameters Model Size Inference Time (ms/img)
(M) (MB) CPU GPU
MultiADS 427.9 1632.64 1824.3 33.7
MultiTypeFCDD 5.5 19.7 69.9 1.7
@@ 6 Conclusion
In this work, we proposed MultiTypeFCDD, a lightweight, unified, and explainable convolutional framework for multi-type anomaly detection. Our key contribution is a weakly supervised model that, using only image-level labels, can simultaneously detect anomalies across multiple object categories and, crucially, differentiate their specific types. This unified approach directly addresses two major gaps in existing literature. First, it eliminates the high overhead of training models per object category. Second, it differentiates between anomaly types, providing the diagnostic specificity that standard anomaly detectors lack. Furthermore, in contrast to massive VLMs that are computationally and memory-intensive, our model is designed as a lightweight and practical alternative, making it suitable for real-world, resource-constrained systems. We demonstrated on the Real-IAD dataset that our framework retains competitive performance against complex state-of-the-art methods while significantly reducing parametric load and inference times. Future work could potentially include extending this framework to adaptively include open-set anomalies, enabling the model to identify and characterise previously unseen anomaly types without explicit retraining.

### v2

@@ 5.2 Multi-Class Anomaly Detection
We also evaluate MultiTypeFCDD in a multi-class anomaly detection setting using a single-channel anomaly output. This allows for fair benchmarking against state-of-the-art multi-class anomaly detection methods, as summarised in Table 5.2. While unsupervised methods show degradation in performance as the proportion of anomalous samples in the training dataset increases, weakly supervised approaches like ours, even with just image-level labels, benefit from additional anomalous samples and show improved performance. MultiTypeFCDD achieves consistently high I-AUROC (above 94%) across different $\alpha$ levels, demonstrating strong anomaly detection capability and generalisation across object categories. Although its AUPRO scores, which indicate anomaly localisation performance, are lower than those of methods such as PatchCore or SoftPatch, this is mainly because these approaches train separate models for each object class, allowing more precise localisation, and the results are averaged. In contrast, unified frameworks like MultiTypeFCDD and UniAD utilise a single model trained across all object classes, which makes fine-grained localisation more challenging but substantially reduces the total number of parameters, requiring only a fraction of the memory and computational overhead while maintaining competitive overall performance.
Table 4: Multi-Class Anomaly Detection Performance for $\alpha$-values 0.1, 0.2, 0.4. The asterisk * indicates that the method uses a single model across all object classes.
Method I-AUROC AUPRO Parameters
0.1 0.2 0.4 0.1 0.2 0.4 (M)
PaDiM [defard2021padim] 81.9 80.1 77.0 86.4 86.5 86.1 68.9
CFlow [gudovskiy2022cflow] 83.9 79.6 78.0 90.7 90.7 90.2 30.4
PatchCore [roth2022towards] 90.4 89.5 88.1 94.2 93.9 92.4 68.9
SoftPatch [jiang2022softpatch] 90.9 90.5 89.3 92.9 92.9 92.5 68.9
SimpleNet [liu2023simplenet] 79.9 75.9 74.7 83.9 83.2 70.5 68.9
DeSTSeg [zhang2023destseg] 85.6 80.3 74.4 86.9 83.2 75.5 37.9
RD [deng2022anomaly] 95.1 94.9 94.7 87.6 83.2 81.5 68.9
UniAD[you2022unified] 84.2 82.8 80.1 87.7 87.3 86.6 7.6
MultiTypeFCDD* 94.1 95.5 96.4 73.3 73.7 78.1 5.5
@@ 5.3 Computational Performance
To emphasise the practical deployability of our approach, we benchmark its computational footprint against MultiADS [sadikaj2025multiads] presented in Table 5. While MultiADS is a potential baseline in multi-type anomaly detection, it employs a VLM trained with image-text prompts, which differs substantially from our approach that uses only images and standard convolutional layers. Consequently, we do not perform a direct comparison under the multi-type anomaly detection setting as it would not meaningfully reflect the strengths of our approach.
Remarkably, MultiTypeFCDD operates with only 5.5 M parameters, marking a significant 98.7% reduction compared to the 427.9 M parameters used in MultiADS. This massive reduction in parametric load directly translates to significantly faster inference times, with ours requiring only 41.7 ms and 1.4ms on CPU and GPU respectively, compared to 1824.3 ms (CPU) and 33.7 ms (GPU) for MultiADS. This comparison demonstrates that complex VLMs are not strictly necessary to achieve reliable multi-type anomaly detection. These findings align with prior observations where large-scale VLMs are noted to be computationally and memory-intensive, restricting their usability in real-time or embedded systems [garcia2019estimation, marshall2025transformers, strubell2020energy]. In contrast, our lightweight architecture achieves competitive performance while substantially reducing memory and inference costs.
@@
Model Parameters Inference Time (ms/img)
(M) CPU GPU
MultiADS [sadikaj2025multiads] 427.9 1824.3 33.7
MultiTypeFCDD 5.5 41.7 1.4
5.4 Ablation Study
Table 6 presents an ablation study on MultiTypeFCDD analysing the impact of backbone, downsampling stage, convolutional blocks, and balanced training on model performance and inference times. We report the mean multi-type anomaly detection metrics across all anomaly classes using the training dataset with $\alpha=0.1$.
Table 6: Ablation study on model performance. We analyse the effects of (a) Backbone, (b) Downsampling Stage, (c) Convolutional Blocks, and (d) Balanced Training Strategy. Default settings are indicated by a checkmark ✓.
Option Params I-AUROC P-AUROC AUPRO Inf. Time (ms/img)
(M) CPU GPU
(a) Backbone
ResNet-18 3.6 94.4 88.7 61.5 11.8 0.4
Inception-v3 4.6 95.5 92.2 69.4 29.6 1.0
Inc-Res-v2 ✓ 5.5 95.7 93.5 75.6 41.7 1.4
(b) Downsampling Stage
Second 3.4 85.9 88.5 45.0 42.4 2.0
Third ✓ 5.5 95.7 93.5 75.6 41.7 1.4
Fourth 34.3 95.6 95.5 70.0 89.6 3.1
(c) Convolutional Blocks
1 3.2 88.5 90.5 71.5 36.4 1.1
2 ✓ 5.5 95.7 93.5 75.6 41.7 1.4
3 7.9 94.0 90.5 71.5 46.3 1.6
(d) Balanced Training
No 5.5 93.5 90.8 73.9 41.7 1.4
Yes ✓ 5.5 95.7 93.5 75.6 41.7 1.4
5.4.1 Backbone:
We compare three pre-trained backbones, ResNet-18, Inception-v3, and Inception-ResNet-v2 (Table 6(a)). While ResNet-18 offers the fastest inference times (11.8 ms CPU, 0.4 ms GPU), Inception-ResNet-v2 achieves the highest detection and localisation metrics (95.7% I-AUROC, 93.5% P-AUROC, 75.6% AUPRO). Despite its higher inference time (41.7 ms CPU, 1.4 ms GPU), the modest parameter increase from just 3.6 M to 5.5 M justifies the substantial performance gains, making it a justifiable backbone choice.
5.4.2 Downsampling Stage:
We compare the performance of the Inception-ResNet-v2 backbone built up to the second, third and fourth downsampling stages (Table 6(b)). With only two stages, performance drops significantly (85.9% I-AUROC, 88.5% P-AUROC, 45.0% AUPRO) due to the insufficient extraction of high-level features. Extending to the fourth stage produces richer features and improves AUROC (95.6% I-AUROC, 95.5% P-AUROC), but causes a substantial increase in parameters (34.3 M) while producing heatmaps that are too small, and resizing such small maps reduces spatial precision and lowers AUPRO to 70.0%. The third-stage backbone balances rich feature extraction and anomaly heatmap reconstruction (95.7% I-AUROC, 93.5% P-AUROC, 75.6% AUPRO), and therefore is the preferred choice.
5.4.3 Convolutional Blocks:
We analyse the effect of varying the number of convolutional blocks appended to the backbone (Table 6(c)). The model underperforms while using only a single block (88.5% I-AUROC, 90.5% P-AUROC, 71.5% AUPRO), indicating insufficient refinement of high-level features. Increasing the number of blocks to two improves all metrics (95.7% I-AUROC, 93.5% P-AUROC, 75.6% AUPRO), demonstrating that an additional capacity is beneficial. However, adding a third block reduces performance (94.0% I-AUROC, 90.5% P-AUROC, 71.5% AUPRO), indicating overfitting due to excess model capacity. Therefore, two convolutional blocks provide the best balance between accuracy, model complexity, and efficiency.
5.4.4 Balanced Training:
We analyse the effect of balanced sampling during training (Table 6(d)). Applying balanced sampling mitigates bias towards more frequent classes such as normal images, stabilises training and improves performance across all metrics (95.7% I-AUROC, 93.5% P-AUROC, 75.6% AUPRO) compared to training without it, thereby justifying its adoption.
@@ 6 Conclusion
In this work, we proposed MultiTypeFCDD, a lightweight, unified, and explainable convolutional framework for multi-type anomaly detection. Our key contribution is a weakly supervised model that, using only image-level labels, can simultaneously detect anomalies across multiple object categories and, crucially, differentiate their specific types. Furthermore, in contrast to massive VLMs that are computationally and memory-intensive, our model is designed as a lightweight and practical alternative, making it suitable for real-world, resource-constrained systems. We demonstrated on the Real-IAD dataset that our framework retains competitive performance against complex state-of-the-art methods while significantly reducing parametric load and inference times.